What is the cost (ins) of heating a hot tub containing The cost of electsoty is 9.00 cents/tkw h). what current (in A) used by the 220 VAC electric heater irthis took os h?
Answer #1

M=1470kg temp change =30C

Heat required to heat water =1470×4186×30=18.46×10^7J

Let energy supplied = Q=75/100Q+heat required to heat water

Q-75/100Q=18.46×10^7

0.925Q=18.46×10^7

Q=19.96×10^7J

1kwh=3.6×10^6

19.96×10^7=55.4kwh

Cost =55.4×9cents=499cents

VIt=55.4

220×I×4.05=55.4

I=0.062A